# Single-shot thermometry of simulated Bose–Einstein condensates using artificial intelligence

### v1

@@ Abstract
Precise determination of thermodynamic parameters in ultracold Bose gases remains challenging due to the destructive nature of conventional measurement techniques and inherent experimental uncertainties. We demonstrate an artificial intelligence approach for rapid, non-destructive estimation of the chemical potential and temperature from single-shot, in situ imaged density profiles of finite-temperature Bose gases. Our convolutional neural network is trained exclusively on quasi-2D ‘pancake’ condensates in harmonic trap configurations. It achieves parameter extraction within fractions of a second. The model also demonstrates zero-shot generalisation across both trap geometry and thermalisation dynamics, successfully estimating thermodynamic parameters for toroidally trapped condensates with errors of only a few nanokelvin despite no prior exposure to such geometries during training, and maintaining predictive accuracy during dynamic thermalisation processes after a relatively brief evolution without explicit training on non-equilibrium states. These results suggest that supervised learning can overcome traditional limitations in ultracold atom thermometry, with extension to broader geometric configurations, temperature ranges, and additional parameters potentially enabling comprehensive real-time analysis of quantum gas experiments. Such capabilities could significantly streamline experimental workflows whilst improving measurement precision across a range of quantum fluid systems.
@@ 1 Introduction
In this paper, we demonstrate a proof-of-principle, non-destructive thermometric technique using artificial intelligence. Our method directly estimates temperature and chemical potential from in situ density profiles, facilitating repeated, rapid measurements on the same atomic sample without significantly perturbing it. Our approach is inspired by recent advances in machine learning to classify and characterise quantum fluid states [18, 19, 20, 21]. We train our artificial intelligence model using simulated density distributions generated from the stochastic Gross–Pitaevskii equation (SGPE), a theoretical framework for modelling finite-temperature dynamics and thermal fluctuations in Bose gases.
We validate our approach across a broad range of temperatures and chemical potentials, and evaluate its estimating capability using unseen density profiles with varied trapping geometries (despite only being trained on harmonically trapped condensates) and during thermalisation (despite only being trained on thermally equilibrated density profiles). Our results demonstrate the robustness and versatility of AI-based thermometry, suggesting its potential as a powerful, precise, and experimentally feasible tool for the investigation of ultracold atomic systems.
@@ 2.2 Dynamics of Bose gases at finite temperatures
We use the stochastic Gross–Pitaevskii equation (SGPE) [24, 25, 26, 27], which statically couples the condensate modes to a thermal bath characterised by chemical potential $\mu$ and temperature $T$ in the grand canonical ensemble. In essence, this is a phenomenologically damped Gross–Pitaevskii equation with additive noise. This approach implements a fluctuation–dissipation theorem that ensures the system relaxes to the correct equilibrium state, without spurious enhancement or depletion of either condensate or thermal components. The SGPE takes the form
@@
$\eta^{*}(\mathbf{x},t)\,\eta(\mathbf{x}^{\prime},t^{\prime})\right>=2\gamma\hbar k_{\mathrm{B}}T\delta(\mathbf{x}-\mathbf{x}^{\prime})\delta(t-t^{\prime});$ (8)
@@ 2.3 Atomic species and trap geometry
kWe consider ultracold, single-species atomic Bose gases of rubidium-87 with scattering length $a_{s}=5.29\,\mathrm{nm}$ [30] and atomic mass $1.44\times 10^{-25}\,\mathrm{kg}$. The temperatures are sampled from $T\sim\mathrm{Uniform}(1,200)\,\mathrm{nK}$, encompassing both deeply degenerate and near-critical regimes. Chemical potentials are sampled from $\mu\sim\mathrm{Uniform}(20,80)\,\mathrm{nK}$ rather than from a regular grid to ensure robust model generalisation. Additional parameters are documented in our open-source code [31].
@@ 3.2 Scaled form of the stochastic Gross–Pitaevskii equation
When generating the training data for the model, we consider a scaled system of units in order to avoid computing very small terms (e.g., of the order of 2superscriptPlanck-constant-over-$\hbar^{2}$ expressed in SI units). In particular, we introduce the following scaling factors: lengths are scaled by $\ell_{x}$ — as defined in section 2.3 — times by $1/\omega_{x}$, energies by $\hbar\omega_{x}$, and temperatures by $\hbar\omega_{x}/k_{\mathrm{B}}$. The condensate field is rescaled as $\tilde{\Phi}(\mathbf{x}_{\mathrm{2D}},t)=\ell_{x}\Phi(\mathbf{x}_{\mathrm{2D}},t)$, and the position and time variables become $\tilde{\mathbf{x}}_{\mathrm{2D}}=\mathbf{x}_{\mathrm{2D}}/\ell_{x}$ and $\tilde{t}=t/\tau$ respectively, where $\tau=1/\omega_{x}$. The Laplacian transforms as ~$\tilde{\nabla}_{\mathrm{2D}}^{2}=\ell_{x}^{2}\nabla^{2}_{\mathrm{2D}}$, while the effective two-dimensional chemical potential and interaction strength are rescaled as $\tilde{\mu}_{\mathrm{2D}}=\mu_{\mathrm{2D}}/\hbar\omega_{x}$ and $\tilde{g}_{\mathrm{2D}}=g_{\mathrm{2D}}/(\hbar\omega_{x}\ell_{x}^{2})$. Neglecting the tildes for convenience, the resulting form of the SGPE is:
@@
$\eta^{*}(\mathbf{x}_{\mathrm{2D}},t)\,\eta(\mathbf{x}_{\mathrm{2D}}^{\prime},t^{\prime})\right>=2\gamma T\delta(\mathbf{x}_{\mathrm{2D}}-\mathbf{x}_{\mathrm{2D}}^{\prime})\delta(t-t^{\prime}).$ (11)
@@
In the following sections, we will first describe the prediction pipeline, starting at layer 4 of the overall neural network, followed by a detailed explanation of the image processing and feature extraction pipeline.
@@ 3.3 Prediction pipeline
which gives estimations for the chemical potential $\mu=y^{6}_{1}$ and temperature $T=y^{6}_{2}$, in nanokelvin. We initialise the weights and biases using the Kaiming scheme [35], as $w^{5}_{jk},\ b^{5}_{j}\sim U(-1/\sqrt{512},1/\sqrt{512})$ and $w^{6}_{jk},\ b^{6}_{j}\sim U(-1/\sqrt{2},1/\sqrt{2})$ (see appendix C in [34] for details).
@@ 3.4 Image processing and feature extraction pipeline
$=\sum_{\sigma=-1}^{1}\sum_{\tau=-1}^{1}\rho(s_{1}+\sigma,\ t_{1}+\tau)\ \mathsf{W}^{1}_{j1}(\sigma,\tau),\quad\text{for }j=1,\dots,12.$
@@
$=\sum_{k=1}^{12}\sum_{\sigma=-1}^{1}\sum_{\tau=-1}^{1}\Xi^{1}_{k}(s_{2}+\sigma,\ t_{2}+\tau)\ \mathsf{W}^{2}_{jk}(\sigma,\tau),\quad\text{for }j=1,\dots,24.$
@@ 3.5.3 Training
Figure 4: Schematic showing the production of the final 48 feature maps from an initial density profile, via the image processing and feature extraction pipeline, which then feed into the prediction pipeline (see also figure 2) to determine estimated values of the chemical potential $\mu$ and temperature $T$. The initial density profile is the same as that of figure 3(a), the (1 of 12) pre-activation image is the same as in figure 3(b), the following post-activation image is the same as in figure 3(c), and the first of the following sample from 12 feature maps is the same as in figure 3(d). Relative to figure 1, this schematic depicts detailed progress, a single channel of each layer at a time, through the image processing and feature extraction pipeline (see section 3.5.2 for details). The end-of-layer outputs from all channels (individual “slices” in figure 1) are inputs to the subsequent layer; the 48 feature maps output by layer 3 are individually globally maximally pooled, as per section 3.3. The complete set of pre-activation images, post-activation images, and feature maps, for each layer, can be seen in in F.
@@ 4.2 Model accuracy
We define ‘accuracy within 5%’ as the proportion of samples with predictions within 5% of their true values — that is, the values which we set as input parameters to the stochastic evolution. Figure 6 illustrates the predictive accuracy, using this metric, for models with 12, 24, and 48 convolutional features, across batch sizes $\beta={16,32,64}$ and $128$. We see that the standard deviation of absolute errors decreases with decreasing batch size, indicating more accurate predictions of both chemical potential and temperature. Smaller batch sizes generally improve generalisation [37] by introducing more stochasticity into the optimisation process, helping the model escape shallow local minima and find solutions that generalise better to unseen data, and this is what we appear to be observing here.
We therefore conclude that reducing the number of features extracted in the convolutional layers to 6, 12, and 24 has minimal impact on predicting the chemical potential [depending on the batch size, the accuracy within 5% remains between 0.98 ($\beta=$ 128) and 1.00 ($\beta=$ 16)] but significantly reduces accuracy for temperature predictions (ranging from 0.73 ($\beta=$ 128) to 0.92 ($\beta=$ 16)). The models with 12, 24, and 48 features have an accuracy within 5% of between 0.99 ($\beta=$ 128) to 1.00 ($\beta=$ 16) for the chemical potential and between 0.88 ($\beta=$ 128) and 0.96 ($\beta=$ 16) for the temperature. Comparing to this model, increasing the features to 16, 32, and 64 does not improve predictions for either variable, as accuracy within 5% remains stable to two decimal places. We have found that the model that extracts 12, 24, and 48 features, with a batch size of 16, offers a good balance between speed and accuracy. All analyses from this point on use this model.
Figure 6: The histograms and corresponding normal distributions demonstrate the absolute error in the model’s prediction of the chemical potential (column 1) and temperature (column 2) to the true values input to our SGPE simulations. The plots share a common $x$-axis for easier comparison. These plots correspond to the models which extract 12, 24 and 48 features.
@@ 4.3 Prediction capability and robustness
In our simulations the bath, with which there is formally both heat and particle exchange, is consistently parametrised throughout by a constant chemical potential, $\mu$, and temperature, $T$. We do not consider dynamics following any form of quench, and the machine learning model is trained only on thermalised Bose gases. As described in equation (9), we use convergence of the atom number as a practical criterion for having achieved thermal equilibrium. However we note that estimated values for $\mu$ and $T$ given by our model appear to trend to the final values more quickly, during the thermalisation process, as shown in figure 7. The model’s ability to estimate parameters during thermalisation, despite training only on equilibrium states, suggests it has learned features that are indicative of the system’s eventual thermodynamic state before reaching equilibrium.
@@
While, as shown in figure 8, the estimated $\mu$ and $T$ trend towards stationary values relatively quickly, we note that the estimated temperature appears in general to be closer to the input value than is the case for the chemical potential. We believe this is because the chemical potential is more associated with the bulk spatial distribution of the condensate, which is very distinct between toroidally and harmonically trapped condensates, whereas the temperature is more associated with small-scale local density fluctuations. This appears also to be supported by our investigations comparing average to maximum pooling.
@@
Both average pooling and maximum pooling are commonly used approaches in, e.g., image recognition. As discussed in D, average pooling gives a general representation of smaller regions in the image whereas maximum pooling emphasises the most prominent features in those regions. It is also possible to think of an equivalency between average pooling, and taking an ensemble average over several noise trajectories; where individual trajectories are stochastic; change associated with averaging over these trajectories is smooth and deterministic. Akin to experiments, however, numerical runs obtained from a single noise realisation hold important physical information; one may interpret a single numerical run of the stochastic Gross–Pitaevskii equation as an independent experimental realisation [24, 41, 42, 43].
@@ 5 Experimental considerations
Whilst our model’s resolution is appropriate for experimental data analysis, real BEC images often include additional complexities such as imaging artefacts, focus variations, and dark counts. Taking such known complexities of the specific imaging system into account would enhance the model’s robustness for experimental use, and should be achievable by augmenting the training data to emulate experimental variations by, for example, applying an appropriate image filter to mimic imperfections in absorption imaging. We also note that, while our model has been trained on in situ imaged position distributions, training a model on an expanding time-of-flight velocity distribution is an in principle straightforward extension, although significantly more demanding to produce the training data if the time-of-flight dynamics are to be taken fully into account.

### v2

# Thermometry of simulated Bose–Einstein condensates using machine learning

@@ Abstract
Precise determination of thermodynamic parameters in ultracold Bose gases remains challenging due to the destructive nature of conventional measurement techniques and inherent experimental uncertainties. We demonstrate a machine learning approach for rapid, non-destructive estimation of the chemical potential and temperature from a single image of an in situ imaged density profiles of finite-temperature Bose gases. Our convolutional neural network is trained exclusively on quasi-2D ‘pancake’ condensates in harmonic trap configurations. It achieves parameter extraction within fractions of a second. The model also demonstrates some zero-shot generalisation across both trap geometry and thermalisation dynamics, successfully estimating the temperature (although not the chemical potential) for toroidally trapped condensates with errors of only a few nanokelvin despite no prior exposure to such geometries during training, and maintaining predictive accuracy during dynamic thermalisation processes after a relatively brief evolution without explicit training on non-equilibrium states. These results suggest that supervised learning can overcome traditional limitations in ultracold atom thermometry, with extension to broader geometric configurations, temperature ranges, and additional parameters potentially enabling comprehensive real-time analysis of quantum gas experiments. Such capabilities could significantly streamline experimental workflows whilst improving measurement precision across a range of quantum fluid systems.
@@ 1 Introduction
In this paper, we demonstrate a proof-of-principle, non-destructive thermometric technique using machine learning Our method directly estimates temperature and chemical potential from in situ density profiles, facilitating repeated, rapid measurements on the same atomic sample without significantly perturbing it. Our approach is inspired by recent advances in machine learning to classify and characterise quantum fluid states [18, 19, 20, 21]. We train our machine learning model using simulated density distributions generated from the stochastic Gross–Pitaevskii equation (SGPE), a theoretical framework for modelling finite-temperature dynamics and thermal fluctuations in Bose gases.
We validate our approach across a broad range of temperatures and chemical potentials, and evaluate its estimating capability using unseen density profiles with varied trapping geometries (despite only being trained on harmonically trapped condensates) and during thermalisation (despite only being trained on thermally equilibrated density profiles). Our results demonstrate the robustness and versatility of machine learning thermometry, suggesting its potential as a powerful, precise, and experimentally feasible tool for the investigation of ultracold atomic systems.
@@ 2.2 Dynamics of Bose gases at finite temperatures
We choose to model the growth of the condensate with the stochastic Gross–Pitaevskii equation (SGPE) [24, 25, 26, 27], which describes a static coupling of the condensate modes to a thermal bath. The growth is towards a state that corresponds to a thermodynamical equilibrium in the grand canonical ensemble defined by our choice of $\mu$ and $T$, which, by definition, must take real values. In essence, the SGPE is a phenomenologically damped Gross–Pitaevskii equation with additive noise. This approach implements a fluctuation–dissipation theorem that ensures the system relaxes to the correct equilibrium state, without spurious enhancement or depletion of either condensate or thermal components. The SGPE takes the form
@@
$\left<\eta^{*}(\mathbf{x},t)\,\eta(\mathbf{x}^{\prime},t^{\prime})\right>=2\gamma\hbar k_{\mathrm{B}}T\delta(\mathbf{x}-\mathbf{x}^{\prime})\delta(t-t^{\prime});$ (8)
@@ 2.3 Atomic species and trap geometry
We consider ultracold, single-species atomic Bose gases of rubidium 87 with scattering length $a_{s}=5.29\,\mathrm{nm}$ [30] and atomic mass $1.44\times 10^{-25}\,\mathrm{kg}$. The temperatures are sampled from $T\sim\mathrm{Uniform}(1,200)\,\mathrm{nK}$, encompassing both deeply degenerate and near-critical regimes. Chemical potentials are sampled from $\mu\sim\mathrm{Uniform}(20,80)\,\mathrm{nK}$ rather than from a regular grid to ensure robust model generalisation. Additional parameters are documented in our open-source code [31].
@@ 3.2 Scaled form of the stochastic Gross–Pitaevskii equation
When generating the training data for the model, we consider a scaled system of units in order to avoid computing very small terms (e.g., of the order of $\hbar^{2}$ expressed in SI units). In particular, we introduce the following scaling factors: lengths are scaled by the harmonic oscillator lengths $\ell_{x,y,z}=\sqrt{{\hbar}/{m\omega_{x,y,z}}}$, times by $1/\omega_{x}$, energies by $\hbar\omega_{x}$, and temperatures by $\hbar\omega_{x}/k_{\mathrm{B}}$. The condensate field is rescaled as $\tilde{\Phi}(\mathbf{x}_{\mathrm{2D}},t)=\ell_{x}\Phi(\mathbf{x}_{\mathrm{2D}},t)$, and the position and time variables become $\tilde{\mathbf{x}}_{\mathrm{2D}}=\mathbf{x}_{\mathrm{2D}}/\ell_{x}$ and $\tilde{t}=t/\tau$ respectively, where $\tau=1/\omega_{x}$. The Laplacian transforms as $\tilde{\nabla}_{\mathrm{2D}}^{2}=\ell_{x}^{2}\nabla^{2}_{\mathrm{2D}}$, while the effective two-dimensional chemical potential and interaction strength are rescaled as $\tilde{\mu}_{\mathrm{2D}}=\mu_{\mathrm{2D}}/\hbar\omega_{x}$ and $\tilde{g}_{\mathrm{2D}}=g_{\mathrm{2D}}/(\hbar\omega_{x}\ell_{x}^{2})$. Neglecting the tildes for convenience, the resulting form of the SGPE is:
@@
$\left<\eta^{*}(\mathbf{x}_{\mathrm{2D}},t)\,\eta(\mathbf{x}_{\mathrm{2D}}^{\prime},t^{\prime})\right>=2\gamma T\delta(\mathbf{x}_{\mathrm{2D}}-\mathbf{x}_{\mathrm{2D}}^{\prime})\delta(t-t^{\prime}).$ (11)
@@
In the following sections, we will first describe the prediction pipeline, starting at layer 4 of the overall neural network, followed by a detailed explanation of the image processing and feature extraction pipeline. We outline the specific mathematical operations used in our convolutional network and contextualise what is often thought of as a ‘black box,’ in a presentation intended to be relevant and accessible for anyone with a physics background.
@@ 3.3 Prediction pipeline
which gives estimations for the chemical potential $\mu=y^{6}_{1}$ and temperature $T=y^{6}_{2}$, in nanokelvin. We initialise the weights and biases using the Kaiming scheme [35], as $w^{5}_{jk},\ b^{5}_{j}\sim U(-1/\sqrt{512},1/\sqrt{512})$ and $w^{6}_{jk},\ b^{6}_{j}\sim U(-1/\sqrt{2},1/\sqrt{2})$ (see appendix C in [34] for details). The final layer does not use an activation function, which is motivated by the physics. The chemical potential can be any real value, and the temperature can be any positive, real value, so no activation function is required (one might, in principle, use the $\mathrm{ReLU}=\max(0,x)$ activation for the temperature, but we did not observe negative temperatures being output from our model). [^3] This contrasts with image classification models, whose final layer typically uses a normalised exponential activation function — commonly referred to as the ‘softmax’ — to produce a probability distribution over output classes [36].
@@ 3.4 Image processing and feature extraction pipeline
[…]
@@
[…]
@@ 3.5.3 Training
Figure 4: Schematic showing the production of the final 48 feature maps from an initial density profile, via the image processing and feature extraction pipeline, which then feed into the prediction pipeline (see also figure 2) to determine estimated values of the chemical potential $\mu$ and temperature $T$. The initial density profile is the same as that of figure 3(a), the (1 of 12) pre-activation image is the same as in figure 3(b), the following post-activation image is the same as in figure 3(c), and the first of the following sample from 12 feature maps is the same as in figure 3(d). Relative to figure 1, this schematic depicts detailed progress, a single channel of each layer at a time, through the image processing and feature extraction pipeline (see section 3.5.2 for details). The end-of-layer outputs from all channels (individual “slices” in figure 1) are inputs to the subsequent layer; the 48 feature maps output by layer 3 are individually globally maximally pooled, as per section $3.$ The complete set of pre-activation images, post-activation images, and feature maps, for each layer, can be seen in in F.
@@ 4.2 Model accuracy
We define ‘accuracy within 5%’ as the proportion of samples with predictions within 5% of their true values — that is, the values which we set as input parameters to the stochastic evolution. Using this metric, figure 6 illustrates the predictive accuracy on the test dataset held back for this purpose, for models with 12, 24, and 48 convolutional features, across batch sizes $\beta={16,32,64}$ and $128$. We see that the standard deviation of absolute errors decreases with decreasing batch size, indicating more accurate predictions of both chemical potential and temperature. Smaller batch sizes generally improve generalisation [38] by introducing more stochasticity into the optimisation process, helping the model escape shallow local minima and find solutions that generalise better to unseen data, and this is what we appear to be observing here.
As summarised in Table 1, reducing the convolutional feature set to 6, 12, and 24 features has negligible effect on chemical-potential estimation but degrades temperature prediction, whereas the 12, 24, and 48 feature configuration gives the best balance of speed and accuracy. Models with 16, 32, and 64 features produced no measurable improvement over the model with 12, 24, and 48 features —accuracies unchanged to two decimal places. We have found that the model that extracts 12, 24, and 48 features, with a batch size of 16, offers a good balance between speed and accuracy. All analyses from this point on use this model.
Table 1: Test dataset accuracy for chemical potential $\mu$ and temperature $T$, defined as the proportion of predictions within $5\%$ of the true values (see § 4.2). Results shown for three model groups (feature sets) across batch sizes $\beta$. The models with the best accuracy within 5% for the chemical potential and temperature are highlighted in bold.
Model (features) Batch size $\beta$ Accuracy within 5% ($\mu$) Accuracy within 5% ($T$)
6, 12, 24 16 1.00 0.92
32 0.99 0.92
64 0.98 0.82
128 0.98 0.73
12, 24, 48 16 1.00 0.96
32 1.00 0.94
64 0.99 0.90
128 0.99 0.88
16, 32, 64 16 1.00 0.96
32 1.00 0.94
64 0.99 0.90
128 0.99 0.88
Figure 6: The histograms and corresponding normal distributions demonstrate the absolute error in the model’s prediction of the chemical potential (column 1) and temperature (column 2) of the held-out test data set to the true values input to our SGPE simulations. The plots share a common $x$-axis for easier comparison. These plots correspond to the models which extract 12, 24 and 48 features.
@@ 4.3 Prediction capability and robustness
4.3.1 Estimation capability during thermalisation
In our simulations the bath, with which there is formally both heat and particle exchange, is consistently parametrised throughout by a constant chemical potential, $\mu$, and temperature, $T$. We do not consider dynamics following any form of quench, and the machine learning model is trained only on thermalised Bose gases. As described in equation (9), we use convergence of the atom number as a practical criterion for having achieved thermal equilibrium. However we note that estimated values for $\mu$ and $T$ given by our model appear to trend to the final values more quickly, during the thermalisation process and after a relatively brief transient, as shown in figure 7. The model’s ability to estimate parameters during thermalisation, despite training only on equilibrium states, suggests it has learned features that are indicative of the system’s eventual thermodynamic state before reaching equilibrium.
@@ 4.3.2 Toroidally trapped condensates
While, as shown in figure 8, the estimated $\mu$ and $T$ trend towards stationary values relatively quickly, the temperature is generally closer to the input value than the chemical potential. This difference reflects that $\mu$ is encoded in the bulk density distribution, which differs qualitatively between harmonic and toroidal traps, whereas $T$ is encoded in small-scale fluctuations that are more geometry-invariant.
Convolutional neural networks exploit weight sharing — each $3\times 3$ filter $\mathsf{W}_{jk}^{\ell}$ is applied identically across all spatial locations, so the same kernel contributes to every receptive field (i.e., each local $3\times 3$ patch of the input). During backpropagation, the gradient of a filter weight is accumulated over all receptive fields, so updating one weight modifies the response of that filter simultaneously at all positions. This enforces a translation-equivariant inductive bias [32], enabling the network to learn recurring local patterns independently of their absolute position. We interpret that convolutional filters scanning local receptive fields extract temperature-linked signatures that remain invariant under a change from harmonic to toroidal confinement.
Two additional factors accentuate the asymmetry: (i) the random transverse anisotropy injected during training (section 2.3) encourages robustness of local features to geometric perturbations, and (ii) the maximally global pooling before the fully connected layers discards absolute position, emphasising distributional summaries of fluctuation statistics rather than any geometry-specific structure.
@@
4.3.3 Ensemble averages
Both average pooling and maximum pooling are commonly used approaches in, e.g., image recognition. As discussed in D, average pooling gives a general representation of smaller regions in the image whereas maximum pooling emphasises the most prominent features in those regions. It is also possible to think of an equivalency between average pooling and taking an ensemble average over several noise trajectories, where individual trajectories are stochastic; change associated with averaging over these trajectories is smooth and deterministic. Akin to experiments, however, numerical runs obtained from a single noise realisation hold important physical information; one may interpret a single numerical run of the stochastic Gross–Pitaevskii equation as an independent experimental realisation [24, 42, 43, 44].
@@ 5.1 Effective pixel size
5.2 Further considerations
[…]
Whilst our model’s resolution is appropriate for experimental data analysis, real BEC images often include additional complexities such as imaging artefacts, focus variations, and dark counts. We also draw attention to very recent independent work, developed in parallel by De Sousa et al. [48], for machine learning thermometry of a dilute, laser cooled (non-quantum-degenerate) vapour of potassium 39 atoms, from experimental data.